# Dynamical scaling study for the estimation of dynamical exponent $z$ of three-dimensional XY spin glass model

### v4

@@ 2.1 Models and Observables for Estimating the Critical Temperature
The relation expressed by Eq. (13) holds only within the context of the NER process in SG systems. Just above the transition temperature, Eq. (13) becomes equivalent to the definition of $z$. This relation was first noted for the Gaussian bond Ising model and gauge glass model[10]. It was later noted for Ising, XY, and Heisenberg SG systems[9]. The crossover between this dynamics and the normal dynamics is observed at the multicritical point in Ising SG systems[2]. However, it is noted that the temperature dependence of $z(T)$ has been confirmed for the SG case, but the validity of this relation in the CG case has not been sufficiently established over the entire temperature range. An analysis based on the aforementioned method has been previously reported[9]. Considering the advances in numerical techniques and available computational power since then, we conducted renewed simulations in the present study.
@@ 2.3 Gaussian Process Regression
In this subsection, we briefly explain the method of dynamic scaling analysis based on Gaussian process regression [11, 12, 6, 7]. This method is applied in all scaling analyses, and it is also used in Sec. 4.1 to assess the validity of the scaling plots. Let us rewrite the scaling relations expressed in Eqs. (10), (14), and (22) in the following form:
@@ 3 Confirmation for the Estimation of the Dynamical Exponent
Figure 2 shows the dynamical correlation function of the 3D FM Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (22) to determine the set of correlation lengths required for an accurate estimation of $z$. In an optimization procedure, we discarded short-range correlations ($r<10$) and data of small $f(r,t)$ values ($f(r,t)<10^{-4}$). The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method. Figure 3 shows the scaling result based on the dynamical scaling law in Eq. (22). The resultant values of correlation length are plotted in Fig. 4 together with those obtained by the second moment method for comparison. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=2.038(2)$. In addition, the uncertainty was estimated as the standard deviation of the bootstrap distribution obtained from 100 resamplings[15]. This value of $z$ is compatible with the value $z=2.0346(3)$ obtained using the latest reliable methods[16], demonstrating the robustness and accuracy of our scaling analysis based on Eq. (22).
@@
Figure 5 shows the dynamical SG correlation function of the 3D $\pm J$ Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (22) to determine the set of correlation lengths required to accurately estimate $z$. During the optimization procedure, short-range correlations ($r<5$) and data with small $f(r,t)$ values ($f(r,t)<5\times 10^{-5}$) were excluded. The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method (Eq. (23)). Figure 6 shows the scaling result obtained using the dynamical scaling law in Eq. (22). The resultant values of correlation length are plotted in Fig. 7. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=5.895(7)$; its uncertainty, defined as the standard deviation of the bootstrap distribution, was assessed using 100 bootstrap trials[15]. This result is consistent with the value $z=5.89(3)$ obtained through extrapolation of the dynamical Binder parameter[2], demonstrating the robustness and accuracy of our scaling analysis approach based on Eq. (22).
@@ 4 Results for 3D $\pm J$ XY model
Figure 8: Size dependence of $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ in the 3D $\pm J$ XY model: (a) $\chi_{\mathrm{SG}}$, where size dependence vanishes for $L\geq 41$, and (b) $\chi_{\mathrm{CG}}$, where size dependence vanishes for $L\geq 51$. The temperatures used for the calculations were $T=0.379$ for $\chi_{\mathrm{SG}}$ and $T=0.4460$ for $\chi_{\mathrm{CG}}$, which are close to the critical point.
@@
To perform this analysis, it is necessary to examine the temperature dependence of $z$ expressed in Eqs. 13 and 17. From the demonstrations described in Secs. 3.1 and 3.2, we note that the reliability of $z$ obtained from the correlation length estimated by the dynamical scaling law has been established. Based on this result, we calculated the SG and CG correlation functions over a temperature range that includes both low- and high-temperature regimes, using $n=256$ and 5 samples. Subsequently, a scaling analysis was conducted according to Eq. (22).
@@
The result is shown in Fig. 9. According to this figure, the temperature dependence of $z$ expressed in Eqs. 13 and 17 is confirmed for both SG and CG susceptibilities. Although it has been reported in other SG systems that this linear temperature dependence breaks down at low temperatures [7], it is considered to remain consistent in the vicinity of the transition temperature. In addition, the intercepts obtained from linear fits to the data do not necessarily vanish; however, the linear fits constructed from the transition temperature and the value of $z$ at the transition point, with the intercept fixed to zero, provide good approximations. These observations indicate that the dynamical scaling analysis based on Eqs. 13 and 17 is required.
@@ 4.1 Dynamical Scaling Analysis
The quantities used in the calculations for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ were $n=256$ and 5 samples. The temperature range used for the calculations is shown in Figs. 10 and 11. For $\chi_{\mathrm{SG}}$, the temperature range was chosen such that the dynamical scaling described by Eq. (10) does not break down and remains valid in the vicinity of the critical temperature. Specifically, including data down to $T\geq 0.410$ yields an apparent critical temperature of $T=0.4143(2)$ in the dynamical scaling analysis, which contradicts the fact that Eqs. (10) are applicable only on the high-temperature side. To avoid this inconsistency while still covering temperatures sufficiently close to the transition, we adopt the range $T\geq 0.420$.
In the scaling analysis for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$, the dynamical exponent $z$ was treated as temperature dependent, following the assumption described in Eqs. (13) and (17). As a result, we obtained $T_{\mathrm{SG}}=0.379(3)$ and $T_{\mathrm{CG}}=0.4460(6)$. In addition, the product of critical exponents $\gamma/z\nu$ and $z\nu$ were estimated as scaling parameters. For $\chi_{\mathrm{SG}}$, $\gamma/z\nu=0.3513(13)$ and $z\nu=8.65(14)$ were obtained; for $\chi_{\mathrm{CG}}$, $\gamma/z\nu=0.3214(6)$ and $z\nu=5.70(1)$ were obtained. These results are compared with those reported in previous studies in Table 1 and Table 2.
@@
Although the authors of some previous studies in which the NER method was used did not provide evidence supporting this theory [4, 9], our analysis of the transition temperature using a similar approach yielded results consistent with it. This discrepancy is attributable to finite-time corrections. The computational MCSs in the present study were approximately $5\times 10^{5}$ MCS, which is longer than the MCSs in previous studies[4, 9]. Figure 12 presents the relaxation data obtained in our study, scaled using the parameters from a previous study [9]. Since the temperature used in the previous study was $T=0.435$, we adopted the temperature range $T\geq 0.460$ for this comparison. The inset shows the distribution of the log-likelihood values obtained via bootstrap sampling, comparing the fits using the parameters from a previous study [9] and those obtained in the present study. To evaluate the log-likelihood functions, we fixed the physical parameters such as $T_{\mathrm{SG}}$, $\gamma/z\nu$, and $z\nu$, and optimized the hyperparameters in the Gaussian process regression for each bootstrap sample. This procedure allows us to evaluate which set of physical parameters is more consistent with the dynamical scaling behavior of the relaxation data. In the prior study, physical quantities were estimated on the basis of relaxation data up to $1\times 10^{5}$ MCS. As shown in Fig. 12, the scaling was successful for data up to $1\times 10^{5}$ MCS. However, when the data were extended to $5\times 10^{5}$ MCS, the scaling failed, which is also clearly reflected in the behavior of the log-likelihood function. This discrepancy is likely attributable to finite-time corrections and supports the validity of our results. Similar issues have been pointed out in finite-size scaling analysis, where the picture can change for small systems[17, 18].
@@
Figure 10: The results of the dynamical scaling analysis for the 3D $\pm J$ XY model based on Eq. (10): (a) relaxation data for $\chi_{\mathrm{SG}}$ and (b) the scaling plot. The scaling plot in (b) was obtained using the parameters $T_{\mathrm{SG}}=0.379(3)$, $\gamma/z\nu=0.3513(13)$, and $z\nu=8.65(14)$.
@@
Figure 11: The results of the dynamical scaling analysis for the 3D $\pm J$ XY model based on Eq. (14): (a) relaxation data for $\chi_{\mathrm{CG}}$ and (b) the scaling plot. The scaling plot in (b) was generated using the parameters $T_{\mathrm{CG}}=0.4460(6)$, $\gamma/z\nu=0.3214(6)$, and $z\nu=5.70(1)$. In this scaling analysis, $z$ is treated as described in Eq. (17).
@@
Figure 12: Scaling plots of the relaxation data obtained in our study, using the parameters from a previous study [9]: $T_{\mathrm{SG}}=0.435$, $\gamma/z\nu=0.368$, and $z\nu=6.25$. The insets show the distribution of the log-likelihood values obtained via bootstrap sampling, comparing the fits using the physical parameters from a previous study [9] and those obtained in the present study. The scaling plot in (a), using data up to $1\times 10^{5}$ MCS, shows a good data collapse. The scaling plot in (b), using data up to $5\times 10^{5}$ MCS, reveals that part of the data in the extended MCS range does not collapse onto the scaling function. The clearly separated distributions of the log-likelihood values show that the two scaling behaviors are distinguishable, which supports the physical conclusion drawn in this study.
@@ 4.2 Estimation of critical exponents
From the inference results and Eq. (1), the value of $z$ is evaluated as $z=6.148(1)$ for the SG case and $z=5.145(4)$ for the CG case. The uncertainties were quantified as the standard deviations of the bootstrap distributions derived from 100 resamplings[15]. These results and the previously mentioned dynamical scaling analysis in Sec. 4.1 can be used to estimate the critical exponents. These results are compared with those of previous studies in Tables 1 and 2.
The obtained results show improved accuracy for both the transition temperature and the critical exponents compared with previously reported findings.
Here, the value of $\gamma/z\nu$ is found to be almost the same as those reported in previous studies employing NER analysis for the SG and CG transition.
From Fig. 8, one can also extract the value of $\gamma/z\nu$ from the slope of $\chi_{\mathrm{SG}}$ at the transition temperature. Using this procedure, we find $\gamma/z\nu=0.315$ for the SG case and $\gamma/z\nu=0.313$ for the CG case. This demonstrates that the deviation from the value obtained by the dynamical scaling analysis is larger in the SG case than in the CG case. These results suggest that the finite-time corrections affect the dynamical-scaling analysis and the susceptibility analysis to different degrees, indicating the necessity of an extrapolation toward the $t\to\infty$ limit in order to obtain more reliable estimates.
In addition, the dynamical exponent $z$ was found to be lower in the CG case than in the SG case, which supports the observation that finite-size effects are more pronounced in the CG sector (Fig. 8).
@@
Figure 14: The scaling plots of the SG and CG correlation functions in the 3D $\pm J$ XY model: (a) analysis of the SG transition with $\eta_{\mathrm{eff}}=0.210(3)$, and (b) analysis of the CG transition with $\eta_{\mathrm{eff}}=0.510(3)$.
@@ 5 Summary and Discussion
We primarily aimed to validate a method for calculating the dynamical exponent from the correlation function, thereby enabling an analysis of SG systems. The obtained dynamical exponent appeared to be reliable, showed good accuracy, and was not limited to SG systems. The results confirm the reliability of this method. Using this approach, we analyzed the 3D $\pm J$ XY model. Compared with conventional methods, our method achieved a higher precision in evaluating the dynamical exponent. Through this analysis, we also examined the temperature dependence of $z(T)$ in the CG transition. In addition, our results support the validity of the spin chirality decoupling picture in SG systems.
Although some previous studies using the NER method did not support the spin chirality decoupling theory[4, 9], our analysis of the transition temperature using a similar method yielded results that support it. This discrepancy is attributable to the finite-time corrections. The computational MCS in this study was approximately $5\times 10^{5}$ MCS, which is longer than the MCSs in previous studies[4, 9], indicating the necessity of considering corrections in the analysis using the dynamical scaling method, which is a future research direction.
The correlations estimated using our method also face problems related to finite-time corrections. Because we do not extrapolate the scaling regime as $t\rightarrow\infty$ for the calculation of the dynamical exponent, our proposed method would suffer a difficult to apply to cases with strong fluctuations. An example of such a situation is the analysis of SG transitions near multi-critical points in the 3D $\pm J$ Ising model[2]. Making the method more widely applicable will require the development of an extrapolation method for $t\rightarrow\infty$. A reliable and systematic extrapolation scheme has recently been proposed in the NER method[16]. We expect that using such a scheme will lead to solutions to such problems. This work is also a future task.

### v5

@@ 2.1 Models and Observables for Estimating the Critical Temperature
The relation expressed by Eq. (13) holds only within the context of the NER process in SG systems. Just above the transition temperature, Eq. (13) becomes equivalent to the definition of $z$. This relation was first noted for the Gaussian bond Ising model and gauge glass model[10]. It was later noted for Ising, XY, and Heisenberg SG systems[9]. The crossover between this dynamics and the normal dynamics is observed at the multicritical point in Ising SG systems[2]. However, while the temperature dependence of $z(T)$ has been widely discussed and generally supported, there remain some differences in the detailed behavior reported in the literature, particularly regarding its quantitative form over the entire temperature range[9, 7]. An analysis based on the aforementioned method has been previously reported[9]. Considering the advances in numerical techniques and available computational power since then, we conducted renewed simulations in the present study.
@@ 2.3 Gaussian Process Regression
In this subsection, we briefly explain the method of dynamic scaling analysis based on Gaussian process regression(GPR) [11, 12, 6, 7]. This method is applied in all scaling analyses, and it is also used in Sec. 4.2 to assess the validity of the scaling plots. Let us rewrite the scaling relations expressed in Eqs. (10), (14), and (22) in the following form:
@@ 3 Confirmation for the Estimation of the Dynamical Exponent
Figure 2 shows the dynamical correlation function of the 3D FM Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (22) to determine the set of correlation lengths required for an accurate estimation of $z$. In an optimization procedure, we discarded short-range correlations ($r<10$) and data of small $f(r,t)$ values ($f(r,t)<10^{-4}$). The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method. Figure 3 shows the scaling result based on the dynamical scaling law in Eq. (22).
The resultant values of the correlation length are plotted in Fig. 4 together with those obtained by the second-moment method for comparison. From the inference results and Eq. (1), we evaluated the dynamic exponent as $z=2.038(2)$, and the corresponding effective exponent as $\eta_{\mathrm{eff}}=0.036(1)$. In addition, the uncertainty was estimated as the standard deviation of the bootstrap distribution obtained from 100 resamplings [15].
This value of $z$ is compatible with the high-precision estimate $z=2.0346(3)$ obtained using the latest reliable methods [16], demonstrating the robustness and accuracy of our scaling analysis based on Eq. (22). The obtained value of $\eta_{\mathrm{eff}}$ is also compatible, within error bars, with those reported in reliable previous studies [14], where $\eta=0.0361(6)$. Although $\eta_{\mathrm{eff}}$ is introduced as an effective exponent in the present scaling framework, the obtained value of $\eta_{\mathrm{eff}}$ shows good agreement with the previous estimate.
@@
Figure 5 shows the dynamical SG correlation function of the 3D $\pm J$ Ising model at the critical temperature. We subsequently conducted a dynamical scaling analysis based on Eq. (22) to determine the set of correlation lengths required to accurately estimate $z$. During the optimization procedure, short-range correlations ($r<5$) and data with small $f(r,t)$ values ($f(r,t)<5\times 10^{-5}$) were excluded. The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method (Eq. (23)). Figure 6 shows the scaling result obtained using the dynamical scaling law in Eq. (22).
The resultant values of correlation length are plotted in Fig. 7. From the inference results and Eq. (1), we evaluated the value of $z$ to be $z=5.895(7)$; its uncertainty, defined as the standard deviation of the bootstrap distribution, was assessed using 100 bootstrap trials[15]. This result is consistent with the value $z=5.89(3)$ obtained through extrapolation of the dynamical Binder parameter[2], demonstrating the robustness and accuracy of our scaling analysis approach based on Eq. (22). The corresponding effective exponent was evaluated as $\eta_{\mathrm{eff}}=-0.183(2)$. This value is compatible, within error bars, with the previous estimate $\eta=-0.22(4)$ reported in previous study [2]. Although $\eta_{\mathrm{eff}}$ is introduced as an effective exponent in the present scaling framework, the obtained value of $\eta_{\mathrm{eff}}$ remains consistent with the earlier result.
@@ 4.1 Size-dependence and the behavior of $z(T)$
Figure 8: Size dependence of $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ in the 3D $\pm J$ XY model: (a) $\chi_{\mathrm{SG}}$, where size dependence vanishes for $L\geq 41$, and (b) $\chi_{\mathrm{CG}}$, where size dependence vanishes for $L\geq 51$. The temperatures used for the calculations were $T=0.4388$ for $\chi_{\mathrm{SG}}$ and $T=0.4533$ for $\chi_{\mathrm{CG}}$, which are close to the critical point.
@@
To perform this analysis, it is necessary to examine the temperature dependence of $z$ expressed in Eqs. 13 and 17. From the demonstrations described in Secs. 3.1 and 3.2, we note that the reliability of $z$ obtained from the correlation length estimated by the dynamical scaling law has been established. Based on this result, we calculated the SG and CG correlation functions over a temperature range that includes both low- and high-temperature regimes. Specifically, the temperatures were chosen in the range $0.300\leq T\leq 0.600$, using $n=256$ and 5 samples. The correlation function was computed up to $5\times 10^{5}$ MCS. Subsequently, a scaling analysis was conducted according to Eq. (22).
@@
The result is shown in Fig. 9. According to this figure, the temperature dependence of $z$ expressed in Eqs. 13 and 17 is confirmed for both SG and CG susceptibilities in the vicinity of the transition temperature. It should be noted that, on the high-temperature side away from the transition temperature, the value of $z$ increases. This behavior arises because, within the time window up to $5\times 10^{5}$ MCS, the correlation length exhibits a tendency toward saturation.
We also note that some previous studies have reported different conclusions regarding the temperature dependence of $z$. For example, it has been reported that $z$ becomes constant on the high-temperature side in the case of the CG transition [9], and that in the Heisenberg SG system $z$ exhibits a linear temperature dependence with a finite intercept [7]. Although both the latter study and the present work were performed in time regions where finite-size effects are sufficiently negligible, we consider that the present results are less affected by correction terms due to the smaller number of degrees of freedom in the system.
@@ 4.2 Dynamical Scaling Analysis
The quantities used in the calculations for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ were $n=256$ and 5 samples, with simulations performed up to $5\times 10^{5}$ MCS. The initial temperature range employed in the simulations was $0.300\leq T\leq 0.600$.
Next, we examine the temperature range of the data used for the dynamical scaling analysis. The dynamical scaling law expressed by Eqs. (10) and (14) may, in principle, hold also on the low-temperature side. However, in the present analysis the scaling fit did not succeed when low-temperature data were included; therefore, in what follows we restrict the analysis to the high-temperature side.
In the SG analysis based on the dynamical scaling law, we take into account the temperature dependence of $z$ assumed in Eq. (13) (and similarly for CG in Eq. (17)). Hence, the temperature range should be chosen so that these assumptions are satisfied. The resulting temperature dependence of $z$ obtained in Sec. 4.1 is shown in Fig. 9. From this figure, an approximately linear temperature dependence of $z$ is observed from around the high-temperature vicinity of the transition toward lower temperatures. Together with the behavior of the correlation length as a function of time at each temperature, we set the upper bound of the temperature range used in the scaling analysis to $T=0.485$ for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$.
We then determine the lower bound of the temperature range by the following procedure. We first set a candidate lower bound $T_{\min}$ for the dataset and perform the GPR-based parameter optimization under the constraint that the estimated transition temperature satisfies $T_{\mathrm{c}}\geq T_{\min}$. If the scaling fit fails (e.g., no satisfactory collapse is obtained, or the optimized $T_{\mathrm{c}}$ hits the lower-bound constraint, $T_{\mathrm{c}}=T_{\min}$), we increase $T_{\min}$ and repeat the same optimization. We finally adopt the smallest $T_{\min}$ for which the scaling fit succeeds stably. As a result, the final temperature ranges are $0.440\leq T\leq 0.485$ for the SG transition and $0.455\leq T\leq 0.485$ for the CG transition.
The corresponding scaling results for SG and CG are shown in Figs. 10 and 11, respectively. From these analyses, the transition temperatures and the scaling parameters, namely the product of critical exponents $\gamma/z\nu$ and $z\nu$, are estimated. Specifically, for $\chi_{\mathrm{SG}}$ we obtain $T_{\mathrm{SG}}=0.4388(3)$, $\gamma/z\nu=0.3528(1)$, and $z\nu=8.67(5)$; for $\chi_{\mathrm{CG}}$ we obtain $T_{\mathrm{CG}}=0.4533(7)$, $\gamma/z\nu=0.3106(4)$, and $z\nu=6.29(10)$.
Here, the value of $\gamma/z\nu$ can also be obtained independently from the slope of the divergence of the susceptibility at the transition temperature. By estimating it from Fig. 8, we obtain $\gamma/z\nu=0.3508\cdots$ for the SG susceptibility and $\gamma/z\nu=0.3093\cdots$ for the CG susceptibility. In general, these values do not necessarily coincide exactly with those obtained from the scaling fit because finite-time corrections can affect the effective slope. Nevertheless, the two estimates are sufficiently close in both cases. This agreement supports the validity of the temperature ranges determined by the above procedure.
@@
Although the authors of some previous studies in which the NER method was used did not provide evidence supporting this theory [4, 9], our analysis of the transition temperature using a similar approach yielded results consistent with it. In the present work, we consider that the primary origin of the discrepancy from the previous estimates is the choice of the temperature range used in the scaling analysis, namely, restricting the dataset to the range where $z(T)$ exhibits an approximately linear temperature dependence. To examine this point more closely, we compared our results with those of Ref. [9] under conditions that mimic the earlier study: we imposed the same lower temperature cutoff ($T\geq 0.460$) and truncated our relaxation data at $1\times 10^{5}$ MCS, which was the maximum simulation time used in Ref. [9].
Specifically, we fixed the physical parameters $(T_{\mathrm{SG}},\gamma/z\nu,z\nu)$ either to the values reported in Ref. [9] or to those obtained in the present study, and for each bootstrap sample we optimized the remaining GPR hyperparameters and compared the resulting log-likelihood values. The results are shown in Fig. 12, where the main panel presents the dynamical scaling plot obtained from our relaxation data truncated at $1\times 10^{5}$ MCS, and the inset shows the distribution of the log-likelihood values over bootstrap samples for the two fixed-parameter fits.
Even under this restriction to $1\times 10^{5}$ MCS, the fits using our parameter set yield, on average, higher log-likelihood values than those using the parameters of Ref. [9]. These observations support the validity of our analysis strategy, in particular the use of only the temperature range in which the assumed linear behavior of $z(T)$ is satisfied.
@@
Figure 10: The results of the dynamical scaling analysis for the 3D $\pm J$ XY model based on Eq. (10): (a) relaxation data for $\chi_{\mathrm{SG}}$ and (b) the scaling plot. The scaling plot in (b) was obtained using the parameters $T_{\mathrm{SG}}=0.4388(3)$, $\gamma/z\nu=0.3528(1)$, and $z\nu=8.67(5)$.
@@
Figure 11: The results of the dynamical scaling analysis for the 3D $\pm J$ XY model based on Eq. (14): (a) relaxation data for $\chi_{\mathrm{CG}}$ and (b) the scaling plot. The scaling plot in (b) was generated using the parameters $T_{\mathrm{CG}}=0.4533(7)$, $\gamma/z\nu=0.3106(4)$, and $z\nu=6.29(10)$. In this scaling analysis, $z$ is treated as described in Eq. (17).
@@
Figure 12: Scaling plots of the relaxation data obtained in our study, using the parameters from a previous study [9]: $T_{\mathrm{SG}}=0.435$, $\gamma/z\nu=0.368$, and $z\nu=6.25$. The insets show the distribution of the log-likelihood values obtained via bootstrap sampling, comparing the fits using the physical parameters from a previous study [9] and those obtained in the present study. The clearly separated distributions of the log-likelihood values show that the two scaling behaviors are distinguishable, which supports the physical conclusion drawn in this study.
@@ 4.3 Estimation of critical exponents
From the inference results and Eq. (1), the value of $z$ is evaluated as $z=5.191(3)$ for the SG case and $z=5.098(4)$ for the CG case. The uncertainties were quantified as the standard deviations of the bootstrap distributions derived from 128 resamplings[15]. These results and the previously mentioned dynamical scaling analysis in Sec. 4.2 can be used to estimate the critical exponents. These results are compared with those of previous studies in Tables 1 and 2.
The obtained results show improved accuracy for both the transition temperature and the critical exponents compared with previously reported findings. In addition, the dynamical exponent $z$ was found to be lower in the CG case than in the SG case, which supports the observation that finite-size effects are more pronounced in the CG sector (Fig. 8).
@@
Figure 14: The scaling plots of the SG and CG correlation functions in the 3D $\pm J$ XY model: (a) analysis of the SG transition with $\eta_{\mathrm{eff}}=0.307(2)$, and (b) analysis of the CG transition with $\eta_{\mathrm{eff}}=0.555(2)$.
@@ 5 Summary and Discussion
We primarily aimed to validate a method for calculating the dynamical exponent from the correlation function, thereby enabling an analysis of SG systems. The obtained dynamical exponent appeared to be reliable, showed good accuracy, and was not limited to SG systems. The results confirm the reliability of this method. Using this approach, we analyzed the 3D $\pm J$ XY model. Compared with conventional methods, our method achieved a higher precision in evaluating the dynamical exponent. Through this analysis, we also examined the temperature dependence of $z(T)$ near both the SG and CG transitions. In addition, our results support the validity of the spin chirality decoupling picture in SG systems.
Although some previous studies using the NER method did not support the spin chirality decoupling theory [4, 9], our analysis of the transition temperature using a similar method yielded results that support it. In the present work, we attribute this difference primarily to how the temperature dependence of the dynamical exponent $z(T)$ is treated in the dynamical scaling analysis: specifically, we restrict the fitting to the temperature range in which $z(T)$ exhibits an approximately linear behavior near the transition, so that the assumed form of $z(T)$ used in the scaling relations is satisfied. This strategy reduces systematic bias arising from applying the $z(T)$ assumption outside its valid range and leads to more reliable estimates of the transition temperatures and scaling parameters.
The dynamical scaling law for the susceptibilities employed in the present analysis [Eqs. (10) and (14)] may, in principle, still be applicable on the low-temperature side, and we cannot categorically rule out such an analysis. However, we confirmed that scaling fits using only the low-temperature data do not succeed in our analysis. On the other hand, it has been reported that a dynamical scaling analysis of the susceptibility as a function of the dynamical correlation length (i.e., susceptibility–correlation-length scaling with the dynamical correlation length used as the time scale) can work even in the SG phase [6, 7]. A possible interpretation is that, in the SG phase, additional nonequilibrium effects such as aging introduce dependencies beyond the observation time $t$, effectively modifying the dynamical scaling form written solely in terms of $t$. Clarifying the origin of this difference is nontrivial and remains an important subject for future work.
The correlations estimated using our method face problems related to finite-time corrections. Because we do not extrapolate the scaling regime as $t\rightarrow\infty$ for the calculation of the dynamical exponent, our proposed method would suffer a difficult to apply to cases with strong fluctuations. An example of such a situation is the analysis of SG transitions near multi-critical points in the 3D $\pm J$ Ising model[2]. Making the method more widely applicable will require the development of an extrapolation method for $t\rightarrow\infty$. A reliable and systematic extrapolation scheme has recently been proposed in the NER method[16]. We expect that using such a scheme will lead to solutions to such problems. This work is also a future task.